OpenStudy (hoothoot23):

An equation is shown below: 8(2x – 14) + 13 = 4x – 27 Part A: Write the steps you will use to solve the equation, and explain each step. (6 points) Part B: What value of x makes the equation true? (4 points

OpenStudy (hoothoot23):

@Aveline

OpenStudy (hoothoot23):

@Michele_Laino

OpenStudy (aveline):

Do you need help with both parts?

OpenStudy (hoothoot23):

yes

rishavraj (rishavraj):

see 8(2x – 14) + 13 = 4x – 27 16x - 112 + 13 = 4x - 27

OpenStudy (aveline):

Ok. 8(2x – 14) + 13 = 4x – 27 Let's start out by using the distributive property. 8(2x – 14) + 13 = 4x – 27 (8)*(2x) + (-14)*(8) + 13 = 4x - 27 16x - 112 +13 = 4x - 27 Now, let's combine the like terms 16x - 99 = 4x - 27 Now we add 99 to each side: 16x - 99 +99 = 4x - 27 +99 16x = 4x +72 And then we subtract 4x from each side 16x - 4x = 4x - 4x +72 12x = 72 Finally, we divide by 12 12x/12 = 72/12 x=6

OpenStudy (hoothoot23):

thank you

OpenStudy (hoothoot23):

what about the second part

OpenStudy (aveline):

That's just the value of x, 6